OpenStudy (anonymous):

What is the equation of the line in standard form that passes through the point (–2, –3) and is parallel to the line y = x + 9? x – y = 1 x + y = 1 x – y = –5 x + y = –5

OpenStudy (anonymous):

k since it is parallel to the line y=x+9 so ots slope must be equal to the slope of y=x+9 slope of y=x+9 is m=1 now using point slope form (-2,-3) y-y0=m(x-x0) y-(-3)=1(x-(-2) simplify it.

OpenStudy (anonymous):

1. Let L be the equation of the line we are looking for. 2. Since L is parallel to the line l: y=x+9, then the slope of l is equal to the slope of L, 3. Since the the slope of l=1, then slope of L=1. 4. Using the point slope formula y-y_0=m(x-x_0) we have y-(-3)=1(x-(-2) => y+3=x+2 => x-y=1

OpenStudy (anonymous):

thank you sooo much!<3